# Guided Speaker Embedding

### v1

@@ III-B Proposed method: Guided speaker embedding
To handle arbitrary numbers of interference speakers, we instead use the joint activity of interference speakers, i.e., non-target speakers’ activity, as shown in the top left of Figure 1. Let be speakers’ speech activities at , where if the -th speaker is active at and otherwise, and be the target speaker’s index. In the proposed method, the input to the encoder is extended from the original acoustic feature as
@@
We also know that the information at frames in which the target speaker is inactive should not be carried to the speaker embedding calculation. We thus zero out the attention weight at the frames by using the target speaker’s activity as a mask and re-normalize the attention weights in the intervals where the target speaker is active:
@@ IV Experimental setup
Our speaker embedding extractor was based on ECAPA-TDNN with 1 024 channels [16]. The model takes 80-dimensional log mel filterbank features extracted using a sliding window of width and shift as input. In the case of the proposed method, target and non-target speech activities were also fed to the model, resulting in 82-dimensional input. The dimensionality of the output speaker embeddings was set to 192. The baseline ECAPA-TDNN was trained using single-speaker utterances cropped to with a mini-batch size of 256. The proposed model was trained using 128 three-speaker partially-overlapped mixtures generated on the fly at each iteration, resulting in 376 samples in a minibatch. Each mixture was generated using utterances cropped to 3–6 seconds with the constraint that the start times of each utterance differ by at least , which is similar to the way used in a multi-speaker ASR study [4]. The mixture ratio was randomly sampled from .

### v2

@@ III-B Proposed method: Guided speaker embedding
To handle arbitrary numbers of interference speakers, we instead use the joint activity of interference speakers, i.e., non-target speakers’ activity, as shown in the top right of Figure 1. Let be speakers’ speech activities at , where if the -th speaker is active at and otherwise, and be the target speaker’s index. In the proposed method, the input to the encoder is extended from the original acoustic feature as
@@
We also know that the information at frames in which the target speaker is inactive should not be carried to the speaker embedding calculation. For the calculation of and to be used for , only frames in which the target speaker is active are used. We also zero out the attention weight at the frames by using the target speaker’s activity as a mask and re-normalize the attention weights in the intervals where the target speaker is active:
@@ IV Experimental setup
Our speaker embedding extractor was based on ECAPA-TDNN with 1 024 channels [16]. The model takes 80-dimensional log mel filterbank features extracted using a sliding window of width and shift as input. In the case of the proposed method, target and non-target speech activities were also fed to the model, resulting in 82-dimensional input. The dimensionality of the output speaker embeddings was set to 192. The baseline ECAPA-TDNN was trained using single-speaker utterances cropped to with a mini-batch size of 256. The proposed model was trained using 128 three-speaker partially-overlapped mixtures generated on the fly at each iteration, resulting in 384 samples in a minibatch. Each mixture was generated using utterances cropped to 3–6 seconds with the constraint that the start times of each utterance differ by at least , which is similar to the way used in a multi-speaker ASR study [4]. The mixture ratio was randomly sampled from .